What would the equation be if the slope is 2/3 and the y-intercept is (9,11)
Answer by unlockmath(1688)

Hello,
We follow the standard form: y=mx+b where m is the slope and b is where the line crosses the y axis.
So plug in the values:
11=2/3(9)+b
Rewritten as:
11=6+b
Subtract 6 from both sides to get:
5=b
There we go. So,
y=2/3(x)+5
Make sense?
RJ
www.math-unlock.com